OpenStudy (april115):

how is the answer 42?

OpenStudy (dumbcow):

great question...i think its 42 because 7*6*i^4 = 42 no but really what was the problem?

OpenStudy (kinggeorge):

42 is the answer to life, the universe, and everything. It's just an inherent part of the universe.

OpenStudy (anonymous):

42 is always the answer in April

OpenStudy (apoorvk):

'42' is the bliss of life!

OpenStudy (april115):

the question was lydia's mold has a rectangular shape that is 7 inches high and measures 2 inches by 3 inches on the bottom.how many cubic inches of space are in lydia's mold? that's the question dumbcow

OpenStudy (asnaseer):

volume = base area times height

OpenStudy (april115):

show me in numbers so i'll understand more please.

OpenStudy (apoorvk):

ohkay, jokes apart here's how: now, volume of a cuboid = product of the three dimensions = length * breadth * height You have the three measurements. multiplying them, you get volume = 42 cubic inches.

OpenStudy (asnaseer):



OpenStudy (asnaseer):

the area of the base is 2 * 3 = 6 square inches

OpenStudy (april115):

thank you very much to both of you.

OpenStudy (asnaseer):

so the volume is 6 * 7 = 42 cubic inches

OpenStudy (april115):

kk got it

OpenStudy (asnaseer):

glad to hear it - strange how you started off the question though - 42 is a well known number in science fiction - see here: http://news.bbc.co.uk/1/hi/7283155.stm

OpenStudy (apoorvk):

@asnaseer Whoa! didn't know that, thanks!

OpenStudy (asnaseer):

yw

OpenStudy (april115):

ummmm.what is 20.0 to the nearest hundredth?

OpenStudy (apoorvk):

That link was cool.

OpenStudy (apoorvk):

20.00

OpenStudy (asnaseer):

april115 - you should post new questions in the list to the left please

OpenStudy (april115):

ok

OpenStudy (asnaseer):

thanks

OpenStudy (apoorvk):

 this one time.